OpenStudy (foxythepirate0984):

Which equation could be used to solve the problem? A cheetah can run 90 feet in 1 second. At that rate, how far can the cheetah run in 30 seconds (d)? A. d = 90 • 30 B. d = 90 ÷ 30 C. d = 30 ÷ 90 D. d = 90 + 30 @youngstudier ╚═( ͡° ͜ʖ ͡°)═╝ ╚═(███)═╝ ╚═(███)═╝ .╚═(███)═╝ ..╚═(███)═╝ …╚═(███)═╝ …╚═(███)═╝ ..╚═(███)═╝ .╚═(███)═╝ ╚═(███)═╝ .╚═(███)═╝ ..╚═(███)═╝ …╚═(███)═╝ …╚═(███)═╝ …..╚(███)╝ ……╚(██)╝ ………(█) ……….*

OpenStudy (mahmood786123):

its simple

OpenStudy (mahmood786123):

the cheetah runs 90 feet in one second

OpenStudy (mahmood786123):

you wanna find out how far the cheetah can run in 30 secondss

OpenStudy (mahmood786123):

90 times 1 = 90

OpenStudy (mahmood786123):

switch the one with a 30

OpenStudy (foxythepirate0984):

So C?

OpenStudy (mahmood786123):

90 times 30 =

OpenStudy (mahmood786123):

=2,700

OpenStudy (mahmood786123):

A

OpenStudy (kendricklamar2014):

\[9 * 1 = 90\] replace the 1 with 30:\[9 * 30 = 270\]

OpenStudy (kendricklamar2014):

Answer: Choice A

OpenStudy (mahmood786123):

the answer is 2,700 not 270

OpenStudy (mahmood786123):

I even checked my answer on a calculator

OpenStudy (mahmood786123):

Hope I helped :)

OpenStudy (kendricklamar2014):

\[9 \times 30 = 270\] @Mahmood786123

OpenStudy (mahmood786123):

check the calculator

OpenStudy (kendricklamar2014):

I did

OpenStudy (mahmood786123):

oops sorry my computer had a glitch there

OpenStudy (kendricklamar2014):

\[9\times300 = 2,700\]

OpenStudy (mahmood786123):

when I entered the 9 times 30 into the calculator it kept showing the answer was 9 times 30

OpenStudy (mahmood786123):

I mean 2,700 not 9 times 30

OpenStudy (mahmood786123):

I just checked again and it said it was 270

OpenStudy (kendricklamar2014):

Yup